The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

         mcnews@astro.uiuc.edu
http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
 http://www.astro.uni-bonn.de/~mcnews/

No. 39 December 12, 1999

Happy holidays!

Contents

News  1
Abstracts of 14 refereed papers 2
4 conference contributions  10
Thesis abstract  12

News

Catalog of Cepheids from the SMC available over the Internet

The OGLE-II microlensing search team announces release of the OGLE Catalog of Cepheids from the Small Magellanic Cloud. The catalog contains 2049 objects and more than 4.7 · 10^{5} BVI measurements of these stars.

The Catalog, finding charts, and BVI individual observations for all objects are available from the OGLE Internet archive:

http://www.astrouw.edu.pl/~ogle or its US mirror
http://www.astro.princeton.edu/~ogle

Andrzej Udalski

Abstracts of Refereed Papers

The Magellanic Stream and the density of coronal gas in the Galactic halo

Chigurupati Murali

CITA, McLennan Labs, University of Toronto, M5S 3H8 Canada Dept. of Astronomy, University of Massachusetts, Amherst 01003, USA

The properties of the Magellanic Stream constrain the density of coronal gas in the distant Galactic halo. We show that motion through ambient gas can strongly heat Stream clouds, driving mass loss and causing evaporation. If the ambient gas density is too high, then evaporation occurs on unreasonably short timescales. Since heating dominates drag, tidal stripping appears to be responsible for producing the Stream. Requiring the survival of the cloud MS IV for 500 Myr sets an upper limit on the halo gas density n_{H} < 10^{−5} cm^{−3} at 50 kpc, roughly a factor of 10 lower than that estimated from the drag model of Moore & Davis (1994). Implications for models of the evolution of gas in galaxy halos are discussed.

Accepted by: ApJ Letters

For preprints, contact: murali@kea.astro.umass.edu
Also available from the URLs: www.astro.umass.edu/∼murali
or http://xxx.lanl.gov/abs/astro-ph/9912168

First Results of a CO Survey of the Large Magellanic Cloud with NANTEN; Giant Molecular Clouds as Formation Sites of Populous Clusters

Yasuo Fukui^{1}, Norikazu Mizuno^{1}, Reiko Yamaguchi^{1}, Akira Mizuno^{1}, Toshikazu Onishi^{1}, Hideo Ogawa^{1,2}, Yoshinori Yonekura^{2}, Akiko Kawamura^{1,3}, Kengo Tachihara^{1}, Kecheng Xiao^{1}, Nobuyuki Yamaguchi^{1}, Atsushi Hara^{1}, Takahiro Hayakawa^{1}, Shigeo Kato^{1}, Rihei Abe^{1}, Hiro Saito^{1}, Satoru Mano^{1}, Ken’ichi Matsunaga^{1}, Yoshihiro Mine^{1}, Yoshiaki Moriguchi^{1}, Hiroko Aoyama^{1}, Shin-ichiro Asayama^{1}, Nao Yoshikawa^{1} and Monica Rubio^{4}

1 Department of Astrophysics, Nagoya University, Chikusa-ku, Nagoya, 464-8602, Japan 2 Department of Earth and Life Sciences, Osaka Prefecture University, Sakai, Osaka 599-8531, Japan 3 Institute of Astronomy, School of Science, The University of Tokyo, Mitaka, Tokyo 181-8588, Japan 4 Departamento de Astronomía, Universidad de Chile, Casilla 36-D, Santiago, Chile

A new survey of the LMC has been completed in 2.6 mm carbon monoxide emission with NAN-TEN. This survey has revealed 107 giant molecular clouds, the first complete sample of giant molecular clouds in a single galaxy at a linear resolution of ∼ 40 pc. The cloud mass ranges from ∼ 6 × 10^{4} to 2 × 10^{6} M_{⊙}, and the total molecular mass has been estimated to be 4–7 × 10^{7} M_{⊙} for a molecular column density of ≥ 1.0 × 10^{21} cm^{−2}, corresponding to 5–10% of the atomic mass. The molecular clouds exhibit a good spatial correlation with the youngest stellar clusters whose ages are ≤10 Myr, demonstrating that cluster formation is on-going in these clouds. On the other hand, they show little correlation with older clusters or with supernova remnants, suggesting that the molecular clouds are being rapidly dissipated in a several Myrs, probably due to the UV photons of massive stars in clusters.

Accepted by: Publ. Astron. Soc. Japan (Vol. 51, No. 6, 1999; NANTEN special issue)

For preprints, contact norikazu@a.phys.nagoya-u.ac.jp
Also available from the URL http://www.a.phys.nagoya-u.ac.jp/NANTEN-PASJ/

HST WFPC2 Imaging of Shocks in Superbubbles

C.-H. Rosie Chen, You-Hua Chu, Robert A. Gruendl, and Sean D. Points

University of Illinois at Urbana-Champaign, USA

Bright X-ray emission has been detected in superbubbles in the Large Magellanic Cloud (LMC), and it is suggested that supernova remnants (SNRs) near the inner shell walls are responsible for this X-ray emission. To identify SNR shocks in superbubble interiors, we have obtained HST WFPC2 emission-line images of the X-ray-bright superbubbles DEM L 152 and DEM L 192 and the X-ray-dim superbubble DEM L 106. We use these images to examine the shell morphology and [S ii]/Hα ratio variations in detail.

Of these three superbubbles, DEM L 152 has the highest X-ray surface brightness, the most filamentary nebular morphology, the largest expansion velocity (∼40 km s^{−1}), and the highest [S ii]/Hα ratio (0.4–0.6). Its [S ii]/Hα ratio increases outwards and peaks in sharp filaments along the periphery. DEM L 192 has a moderate X-ray surface brightness, a complex but not filamentary morphology, a moderate expansion velocity (35 km s^{−1}), and a low [S ii]/Hα ratio (∼0.15). DEM L 106 is not detected in X-rays. Its shell structure is amorphous and has embedded dusty features; its expansion velocity is <10 km s^{−1}.

None of the three superbubbles show morphological features in the shell interior that can be identified as directly associated with SNR shocks, indicating that the SNR shocks have not encountered very dense material. We find that the [S ii]/Hα ratios of X-ray-bright superbubbles are strongly dependent on the UV radiation field of the encompassed OB associations. Therefore, a tight correlation between [S ii]/Hα ratio and X-ray surface brightness in superbubbles should not exist. We also find that the filamentary morphologies of superbubbles are associated with large expansion velocities and bright X-ray emission.

Accepted by: The Astronomical Journal

For preprints, contact c-chen@astro.uiuc.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9912150
or by anonymous ftp at ftp.astro.uiuc.edu, cd /pub/c-chen/

The Low End of the Initial Mass Function in Young LMC Clusters: I. The Case of R136

M. Sirianni^{1,2}, A. Nota^{2,4}, C. Leitherer^{2}, G. De Marchi^{3}, and M. Clampin^{2}

1 Department of Physics and Astronomy, Johns Hopkins University, 3400 N. Charles St., Baltimore, MD, 21218, USA 2 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 3 European Southern Observatory, Karl-Schwarzschild-Str.2, 85748 Garching, Germany 4 Affiliate with the Astrophysical Division, Space Science Department of the European Space Agency

We report the result of a study in which we have used very deep broadband V and I WFPC2 images of the R136 cluster in the Large Magellanic Cloud from the HST archive, to sample the luminosity function below the detection limit of 2.8 M_{⊙} previously reached. In these new deeper images, we detect stars down to a limiting magnitude of m_{F 555W} = 24.7 (≃ 1 magnitude deeper than previous works), and identify a population of red stars evenly distributed in the surrounding of the R136 cluster. A comparison of our color-magnitude diagram with recently computed evolutionary tracks indicates that these red objects are pre-main sequence stars in the mass range 0.6 – 3 M_{⊙}. We construct the initial mass function (IMF) in the 1.35 – 6.5 M_{⊙} range and find that, after correcting for incompleteness, the IMF shows a definite flattening below ≃ 2 M_{⊙}. We discuss the implications of this result for the R136 cluster and for our understanding of starburst galaxies formation and evolution in general.

Accepted by: The Astrophysical Journal

For preprints, contact sirianni@pha.jhu.edu
Also available from the URL http://adcam.pha.jhu.edu/~sirianni/r136/r136prep.html

Chemical Abundances from Magellanic Cloud B Stars

A.J. Korn^{1,2}, S.R. Becker^{1}, C.A. Gummersbach^{2} and B. Wolf^{2}

1 Universitäts-Sternwarte München (USM), Scheinerstr. 1, D-81679 München, Germany 2 Landessternwarte (LSW), Königstuhl, D-69117 Heidelberg, Germany

In recent years, B stars have been established as reliable tracers of present day abundance patterns. They yield an independent set of abundances from He to Fe which can be cross-checked against other sources such as H ii regions and cool supergiants. In this paper, we present purely spectroscopic analyses of 9 non-supergiant B stars (13 < m_{V} < 15), 5 of which located in the young populous clusters NGC 1818, 2004 (LMC) and NGC 330 (SMC), the other 4 being MC field stars. CASPEC spectra (4000 Å < λ < 5000 Å) with R ∼ 20 000 and S/N ∼ 100 are used to determine T_{eff}, log g, ξ and the abundances of oxygen and silicon simultaneously. Particular attention is paid to a consistent treatment of the metallicity of the underlying atmosphere. Kurucz ATLAS 9 LTE atmospheres and DETAIL/SURFACE non-LTE line formation (H, He, C, N, O, Mg, Al, Si and Fe) are used throughout the analysis.

While we find little difference between the abundances from cluster members and field stars, the behaviour of N is surprising: it is generally enriched with respect to the H ii value by 0.6 to 1.0 dex. In the case of NGC 1818/D1 (the only spectroscopic main sequence star in our sample), this enrichment is coupled with a depletion of C and an enrichment of He, strongly suggesting that these abundances have already been altered through some kind of mixing with CNO-processed material.

Accepted by: Astronomy & Astrophysics

For preprints, contact ajkorn@usm.uni-muenchen.de
Also available from the URL www.usm.uni-muenchen.de/people/korn/Bstar.html

R4 and its circumstellar nebula: evidence for a binary merger?

A. Pasquali^{1}, A. Nota^{2}, N. Langer^{3}, R.E. Schulte-Ladbeck^{4}, M. Clampin^{2}

1 ESO-ST/ECF, Karl-Schwarzschild-Strasse 2, D-85748 Garching bei München, Germany 2 STScI, 3700 San Martin Drive, Baltimore, MD 21218, USA 3 Institut für Physik, Universität Potsdam, Postfach 601553, D-14415 Potsdam, Germany 4 University of Pittsburgh, Pittsburgh, PA 15260, USA

We present new, NTT longslit spectroscopy of the B[e] supergiant in the binary system R4 in the Small Magellanic Cloud. The data show extended, forbidden N and S emissions which are typical signatures of circumstellar matter. Their extension along the space axis of the slit defines an angular size of 8.6′′ which translates into a linear size of 2.4 pc. The N emission lines also show the velocity structure of a bipolar outflow expanding at 100 km s^{−1} on average. This implies that, for a measured radius of 1.2 pc, the outflow originated about 1.2 × 10^{4} yr ago. The line flux ratio [NII]6584/[SII]6717 indicates that the nebula is nitrogen enriched and therefore it has been ejected from the central star.

This is the first bipolar, ejection nebula detected around a well-established B[e] supergiant. The bipolar morphology and the chemical enrichment shown by the nebula associated with R4 are consistent with the picture of a binary merger (Langer & Heger 1998), in which R4 was originally a system composed by a close pair and a third star (the observed A companion). The close pair merged into a single star and the merging process produced a circumstellar nebula that was later shaped by the ensuing B star wind.

Since the bipolar morphology, the kinematics and the enriched chemical composition make the nebula surrounding R4 very similar to the observed LBV nebulae, our findings imply that at least a few LBV outbursts and nebulae might well be the result of the merging process of two massive stars.

Accepted by: The Astronomical Journal For preprints, contact apasqual@eso.org

or by anonymous ftp at ftp.hq.eso.org, cd /users/apasqual/outgoing

An X-Ray Spectroscopic Study of the SMC X-1/Sk 160 System

Patrick S. Wojdowski^{1,3}, George W. Clark^{1}, and Timothy R. Kallman^{2}

1 Massachusetts Institute of Technology, Cambridge, MA, USA 2 NASA/Goddard Space Flight Center, Greenbelt, MD, USA 3 Current Address: Lawrence Livermore National Laboratory, USA

We have investigated the composition and distribution of the wind of Sk 160, the supergiant companion of the X-ray star SMC X-1, by comparing an X-ray spectrum of the source, obtained with the ASCA observatory, during an eclipse with the computed spectra of reprocessed radiation from circumstellar matter with various density distributions. We show that the metal abundance in the wind of Sk 160 is no greater than a few tenths of solar, as has been determined for other objects in the Magellanic Clouds. We also show that the observed X-ray spectrum is not consistent with the density distributions of circumstellar matter of the spherically symmetric form derived for line-driven winds, nor with the density distribution derived from a hydrodynamic simulation of the X-ray perturbed and line-driven wind by Blondin & Woo (1995).

Accepted by: The Astrophysical Journal

For preprints, contact patrickw@virgo.llnl.gov
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9912035

The DENIS Point Source Catalogue towards the Magellanic Clouds

M-R. Cioni^{1}, C. Loup^{2}, H.J. Habing^{1}, P. Fouqué^{6,3}, E. Bertin^{2}, E. Deul^{1}, D. Egret^{14}, C. Alard^{15}, B. de Batz^{3}, J. Borsenberger^{2}, M. Dennefeld^{2}, N. Epchtein^{4}, T. Forveille^{5}, F. Garzón^{7}, J. Hron^{8}, S. Kimeswenger^{9}, F. Lacombe^{3}, T. Le Bertre^{15}, G. A. Mamon^{2,3}, A. Omont^{2}, G. Paturel^{10}, P. Persi^{11}, A. Robin^{12}, D. Rouan^{3}, G. Simon^{15}, D. Tiphène^{3}, I.

Vauglin^{10} and S. Wagner^{13}

1 Leiden Observatory, Postbus 9513, NL–2300 RA Leiden, The Netherlands 2 Institut d’Astrophysique de Paris (CNRS UPR 341), 98 bis Bd. Arago, F–75014 Paris, France 3 Observatoire de Paris, 5 place J. Janssen, F–92195 Meudon Cedex, France 4 Observatoire de la Côte d’Azur, Département Fresnel, F–06304 Nice Cedex 04, France 5 Observatoire de Grenoble, 414 rue de la Piscine, Domaine Universitaire de Saint Martin d’Hères, F-38041 Grenoble, France 6 European Southern Observatory, Casilla 19001, Santiago 19, Chile 7 Instituto de Astrofisica de Canarias E–38200 La Laguna, Tenerife, Spain 8 Institut für Astronomie der Universität Wien, Turkenschanzstrasse 17, A–1180 Wien, Austria 9 Institut für Astronomie, Innsbruck University A–6020 Innsbruck, Austria 10 CRAL–Observatoire de Lyon F–69561 Saint–Genis Laval Cedex, France 11 Istituto di Astrofisica Spaziale, Area di Ricerca Roma–Tor–Vergata, I–00044 Italy 12 Observatoire de Besançon BP 1615, F-25010 Besançon Cedex, France 13 Landessternwarte Heidelberg, Königstuhl, D–69117 Heidelberg, Germany 14 CDS, Observatoire Astronomique de Strasbourg (CNRS UMR 7550), 11 rue de l’Université, F–67000 Strasbourg, France 15 Observatoire de Paris, 65” Avenue de l’Observatoire, F–75014 Paris, France

We have compiled the first version of a complete near infrared Point Source Catalogue (PSC) towards the Magellanic Clouds (MCs) extracted from the data obtained with the Deep Near Infrared Survey of the Southern Sky – DENIS (Epchtein et al. 1997). The catalogue covers an area of of 19.87 × 16 square degrees centered on (α, δ) = (5^{h}27^{m}20^{s},−69^{o}00′00′′) for the Large Magellanic Cloud (LMC) and 14.7 × 10 square degrees centered on (α, δ) = (1^{h}02^{m}40^{s},−73^{o}00′00′′) for the Small Magellanic Cloud (SMC). It contains 1, 314, 009 sources towards the LMC and 318, 631 sources towards the SMC each detected in at least 2 of the 3 photometric bands involved in the survey (I, J, K_{s}). 70% of the detected sources are real members of the Magellanic Clouds, respectively and consist mainly of red giants, asymptotic giant branch stars and super-giants. The observations have all been made with the same instrument and therefore provide a homogeneous set of photometric data.

Submitted to: Astronomy & Astrophysics

For preprints, contact mrcioni@strw.leidenuniv.nl

Mass-loss rates and dust-to-gas ratios for obscured Asymptotic Giant Branch stars of different metallicities

Jacco Th. van Loon

Institute of Astronomy, Madingley Road, Cambridge CB3 0HA, United Kingdom

The mass-loss rates and dust-to-gas ratios of obscured Asymptotic Giant Branch (AGB) stars are investigated for samples with different initial metallicities: in the Small and Large Magellanic Clouds (SMC & LMC) and in the Milky Way. The properties of their circumstellar envelopes can be explained in a consistent way if, both for obscured M-type AGB stars and for obscured carbon stars, the total (gas+dust) mass-loss rate depends only weakly on initial metallicity whilst the dust-to-gas ratio depends approximately linearly on initial metallicity.

Accepted by: Astronomy & Astrophysics Main Journal For preprints, contact jacco@ast.cam.ac.uk

The Optical Gravitational Lensing Experiment. Cepheids in the Magellanic Clouds. V. Catalog of Cepheids from the Small Magellanic Cloud

A. Udalski^{1}, I. Soszyński^{1}, M. Szymański^{1}, M. Kubiak^{1}, G. Pietrzyński^{1}, P. Woźniak^{2}, and K. Żebruń^{1}

1 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2 Princeton University Observatory, Princeton, NJ 08544-1001, USA

We present the Catalog of Cepheids from the SMC which contains data for 2049 objects detected in the 2.4 square degree area of central parts of the SMC. For each object period, BVI photometry, astrometry, and R_{21}, ϕ_{21} parameters of the Fourier decomposition of I-band light curve are provided. The Catalog is based on observations collected during the OGLE-II microlensing survey.

Tests of completeness performed in overlapping parts of adjacent fields indicate that completeness of the Catalog is very high: ≈ 92%. Statistics and distributions of basic parameters of Cepheids are also presented.

All presented data, including individual BVI observations (≈ 4.7 · 10^{5} BVI measurements), are available from the OGLE Internet archive.

Submitted to: Acta Astronomica Vol. 49

For preprints, contact udalski@astrouw.edu.pl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9912096

The Optical Gravitational Lensing Experiment. Cepheids in Star Clusters from the Magellanic Clouds

G. Pietrzyński and A. Udalski

Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland

We present Cepheids located in the close neighborhood of star clusters from the Magellanic Clouds. 204 and 132 such stars were found in the LMC and SMC, respectively. The lists of objects were constructed based on catalogs of Cepheids and star clusters, recently published by the OGLE-II collaboration. Location of selected Cepheids on the sky indicates that many of them are very likely cluster members. Photometric data of Cepheids and clusters are available from the OGLE Internet archive.

Submitted to: Acta Astronomica Vol. 49

For preprints, contact pietrzyn@astrouw.edu.pl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9912190

The Optical Gravitational Lensing Experiment. Catalog of Star Clusters from the Large Magellanic Cloud

G. Pietrzyński^{1}, A. Udalski^{1}, M. Kubiak^{1}, M. Szymański^{1}, P. Woźniak^{2}, and K. Żebruń^{1}

1 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2 Princeton University Observatory, Princeton, NJ 08544-1001, USA

We present the catalog of star clusters found in the area of about 5.8 square degree in the central regions of the Large Magellanic Cloud. It contains data for 745 clusters. 126 of them are new objects. For each cluster equatorial coordinates, radius, approximate number of members and cross-identification are provided. Photometric data for all clusters presented in the catalog and Atlas consisting of finding charts and color-magnitude diagrams are available electronically from the OGLE Internet archive.

Submitted to: Acta Astronomica Vol. 49

For preprints, contact pietrzyn@astrouw.edu.pl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9912187

Updating the census of star clusters in the Small Magellanic Cloud.

E. Bica^{1} and C. M. Dutra^{1}

1 Universidade Federal do Rio Grande do Sul, Instituto de Fisica, CP 15051, Porto Alegre 91501-970, RS, Brazil

Surveys using CCD detectors are retrieving bright and faint cataloged clusters and revealing new ones in the Magellanic Clouds. This paper discusses the contribution of the OGLE Survey to the overall census of star clusters in the SMC. A detailed cross-identification indicates that the new objects in the SMC OGLE catalog are 46. The increase in the number of cataloged clusters is ≈ 7%, the total sample being ≈ 700. This updated census includes embedded clusters in H ii regions and a density range attaining loose systems.

Accepted by: The Astronomical Journal (tentatively scheduled for the March 2000 issue)

Also available from the URL http://www.if.ufrgs.br/~dutra/catpage.html (incl. one electronic

table)

The Tip of the Red Giant Branch Distance to the Large Magellanic Cloud

S. Sakai^{1}, D. Zaritsky^{2}, and R. Kennicutt^{2}

1 KPNO/NOAO, P.O.Box 26732, Tucson AZ 85726, USA 2 Steward Observatory, University of Arizona, Tucson, AZ, USA

We present the I–band luminosity function of the red giant branch stars in the Large Magellanic Cloud (LMC) using the data from the Magellanic Clouds Photometric Survey (Zaritsky, Harris & Thompson, 1997). Selecting stars in uncrowded, low–extinction regions, a discontinuity in the luminosity function is observed at I_{0} = 14.54 mag. Identifying this feature with the tip of the red giant branch (TRGB), and adopting an absolute TRGB magnitude of −4.05 ± 0.04 mag based on the calibration of Lee, Freedman & Madore (1993), we obtain a distance modulus of 18.59 ± 0.09 (random) ±0.16 matisse. We present the I-band luminosity function of the red giant branch stars in the Large Magellanic Cloud (LMC) using the data from the Magellanic Clouds Photometric Survey (Zaritsky, Harris & Thompson, 1997). Selecting stars in uncrowded, low-extinction regions, a discontinuity in the luminosity function is observed at I_{0} = 14.54 mag. Identifying this feature with the tip of the red giant branch (TRGB), and adopting an absolute TRGB magnitude of −4.05 ± 0.04 mag based on the calibration of Lee, Freedman & Madore (1993), we obtain a distance modulus of 18.59±0.09 (random) ±0.16 (systematic) mag. If the theoretical TRGB calibration provided by Cassisi & Salaris (1997) is adopted instead, the derived distance would be 4larger by 0.09 mag (1-sigma) than the value that is most commonly used in the extragalactic distance scale calibrated by the period-luminosity relation of the Cepheid variable stars. Our TRGB distance modulus agrees with several RR Lyrae distances to the LMC based on HIPPARCOS parallaxes. Finally, we note that using the same MCPS data, we obtain a distance modulus of 18.29 ± 0.03 mag using the red clump method, which is shorter by 0.3 mag compared to the TRGB estimate.

Accepted by: The Astronomical Journal For preprints, contact shoko@noao.edu

Abstracts of Non-Refereed Papers

Dust Enshrouded AGB Variables in the LMC

Patricia Whitelock^{1} and Michael Feast^{2}

1 SAAO, P O Box 9, Observatory, 7935, South Africa 2 University of Cape Town, South Africa

The luminosities and periods of obscured AGB stars in the LMC are discussed using a combination of ISO and ground-based infrared photometry. The bolometric luminosities of these stars fall close to an extrapolation of the period-luminosity relation derived for oxygen-rich Mira variables, with both oxygen- and carbon-rich stars falling close to the same relation. It has been known for many years that there are, in the Magellanic Clouds, significant numbers of large amplitude variables which have considerably higher luminosities than the period-luminosity relation would predict. Many of these can be shown to be undergoing hot bottom burning. It is speculated that all large amplitude AGB stars with luminosities significantly higher than indicated by the period-luminosity relation are undergoing hot bottom burning.

To appear in: Workshop on “The Changes in Abundances in AGB Stars”, Mem. Soc. Astron. Italy For preprints, contact paw@saao.ac.za

Also available from http://xxx.lanl.gov/abs/astro-ph/9911424

The bulge globular cluster NGC 6553: Observations with HST’s WFPC2, STIS and NICMOS

S.F. Beaulieu^{1}, G.F. Gilmore^{1}, R.A. Johnson^{1}, S.J. Smartt^{1}, N. Tanvir^{2}, and B. Santiago^{3}

1 Institute of Astronomy, Cambridge, UK 2 Department of Physical Science, University of Hertfordshire, UK 3 Universidade Federal do Rio Grande do Sul, Porto Alegre, Brazil

As part of a large HST project to study the formation and evolution of rich star clusters in the Large Magellanic Cloud, we present results for the Galactic bulge metal-rich globular cluster NGC6553. HST observations using WFPC2, NICMOS and STIS were obtained for this cluster. The primary reason for studying NGC6553 is to transform our NICMOS and STIS magnitudes directly into absolute magnitudes. This is particularly important for determining the low mass end of the IMF. NGC 6553 was chosen because its metallicity, [Fe/H] ≈ −0.2, is representative of the metallicities of the young and intermediate age LMC clusters.

To appear in: Massive Stellar Clusters, ASP Conf. Series

For preprints, contact beaulieu@ast.cam.ac.uk

The First Five Minutes of a Core Collapse Supernova: Multidimensional Hydrodynamic Models

K. Kifonidis^{1}, T. Plewa^{2,1}, H.-Th. Janka^{1} and E. Müller^{1}

1 Max-Planck-Institut für Astrophysik, D-85740 Garching, Germany 2 Nicolaus Copernicus Astronomical Center, Bartycka 18, 00716 Warsaw, Poland

We present results of high-resolution two-dimensional simulations which follow the first five minutes of a core collapse supernova explosion in a 15 solar mass blue supergiant progenitor. The computations start shortly after core bounce and include neutrino-matter interactions by using a light-bulb approximation for the neutrinos, and a treatment of the nucleosynthesis due to explosive silicon and oxygen burning. We find that newly formed iron-group elements are distributed throughout a significant fraction of the stellar helium core by the concerted action of convective and Rayleigh-Taylor instabilities. Fast moving nickel mushrooms with velocities up to 4000 km/s are observed. A continuation of the calculations to later times, however, indicates, that the iron velocities observed in SN 1987 A cannot be reproduced due to a strong deceleration of the clumps during their interaction with the dense shell left behind by the shock at the He/H interface. Therefore, we cannot confirm the claim that convective ”premixing” of the nickel in the early phases of the explosion solves the problem of the high iron velocities.

To appear in: R. Diehl, ed., Proceedings of the Workshop on “Astronomy with Radioactivities” held at Ringberg Castle, Tegernsee, Germany, Sep. 29th - Oct. 2nd, 1999, Max-Planck-Institut für Extraterrestrische Physik, Garching

For preprints, contact kok@mpa-garching.mpg.de

The Distance to the Large Magellanic Cloud

Brad K. Gibson^{1}

1 Center for Astrophysics & Space Astronomy, University of Colorado, Boulder, CO 80309-0389

A summary of recent Large Magellanic Cloud distance determinations μ_{LMC} is presented, with an eye towards pinpointing the source(s) of the resulting large discrepancies encountered between some of the techniques. Thirty-eight recent (1998-1999) measurements of μ_{LMC} are highlighted, the extrema for which (18.07 versus 18.74) are inconsistent with one another at the 3σ level. The lack of overlap between the results of many of the techniques, simply employing the authors’ quoted uncertainties, is a clear indication that unrecognized uncertainties, both random and systematic, plague many of the published results. While μ_{LMC} almost certainly lies within ∼13% of 18.5 (i.e., between 18.20 and 18.75), to those of us outside the LMC “community”, no single compelling argument has been put forth that reconciles the wildly disparate values presented thus far. A ∼13% uncertainty in the LMC distance corresponds to a ∼13% uncertainty in the Cepheid-based extragalactic distance scale.

To appear in: Journal of the Italian Astronomical Society (MemSAI)

For preprints, contact bgibson@casa.colorado.edu
Also available from the URL http://casa.colorado.edu/∼bgibson

Thesis Abstract

The complex nature of the ISM in the SMC: an H i and infrared study

Snežana Stanimirović

University of Western Sydney Nepean (UWSN), P.O. Box 10, Kingswood, NSW 2747, Australia Australia Telescope National Facility (ATNF), CSIRO, P.O. Box 76, Epping, NSW 2121, Australia Current address: Arecibo Observatory, HC 3 Box 53995, Arecibo, Puerto Rico 00612

This thesis presents the results of a combination of new observations with the Parkes telescope of neutral hydrogen (H i) in the Small Magellanic Cloud (SMC) with an Australia Telescope Compact Array (ATCA) aperture synthesis mosaic. The new data consists of a set of images sensitive to all angular (spatial) scales between 98 arcsec (30 pc) and 4^{∘} (4 kpc). The data are used to study the H i distribution and mass, the velocity field and rotation curve of the SMC, as well as to probe the 3-D structure of the SMC. A kinematic study of the H i data reveals the existence of three supergiant shells which were previously undetectable in the ATCA data alone. The H i spatial power spectrum has been investigated over a range of contiguous scale sizes wider than those previously achieved in any other galaxy, including our own. It closely obeys the relation, P (k) ∝ k^{γ}, suggesting an alternative model for the interstellar medium (ISM) of the SMC, having fractal nature and consisting of a hierarchy of H i cloud structures.

This thesis also demonstrates that the infrared data obtained with the Infrared Astronomical Satellite (IRAS) for the SMC can be successfully reconstructed with much higher resolution using the Pyramid Maximum Entropy algorithm. The new infrared (IR) data are used to study the integrated IR spectrum, the dust temperature and dust column density in the SMC. The high resolution H i and IR data enable an investigation of the spatial correlation of dust and gas and the assumption of the dust and gas being well-mixed in the ISM. The spatial power spectrum of the dust column density shows that, as with the H i power spectrum, there is no preferred scale size for dust clouds. The remarkable similarity of the spatial power spectra for the H i and dust column density distributions suggests a unique and ubiquitous hierarchical structure organisation for the ISM in the SMC. Such an organisation is likely to be governed by the Kolmogorov type turbulence and could be described by the fractal nature with the volume fractal dimension of 2.2.

Ph.D. Thesis accepted by the University of Western Sydney Nepean, Australia, on November, 29^{th} 1999, under the direction of: Assoc. Prof. Graeme White (UWSN), Dr. Paul Jones (UWSN), Dr. Lister Staveley-Smith (ATNF) and Dr. Raymond Haynes (ATNF). For copies, contact Snežana Stanimirović: sstanimi@naic.edu